If a variable has a distribution that is​ bell-shaped with mean

28

and standard deviation

7​,

then according to the Empirical​ Rule,

68.0%

of the data will lie between which​ values?
Answer #1

Solution:

Given in the question

Mean = 28

Standard deviation = 7

According to the empirical rule

68% of the data will lie between +/- 1 standard deviation from the mean

So lower value = mean - standard deviation = 28 - 7 = 21

Upper value = mean + standard deviation = 28 + 7 = 35

So 68% of the data will lie between 21 to 35.